OpenStudy (anonymous):

solve for x 1/ 36 = 6^(x–3) solve for x 1/ 36 = 6^(x–3) @Mathematics

OpenStudy (jamesj):

\[ 1/36 = 1/(6^2) = 6^{-2} \] Therefore if 1/ 36 = 6^(x–3), then \[ 6^{-2} = 6^{x-3} \] This is true if and only if -2 = x - 3 and therefore ....

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

\[6^{-2} = 6^{x-3} => -2 = x-3 => x = 1\]